Director's Time

Observers: B.Rothberg, T. Rochais (mountain), O. Kuhn, M. (Tucson)

ISA: B.Rothberg
Telescope Operator: D.Gonzalez-Huerta (LBTO)
AO Support: J.Christou
Telescope Support: K.Newton (LBTO)
Software Support: C.Biddick
Instrument Support: J.Morris (LBTO)
Engineering Support: M.Gusick

Summary

First Half E-time, 2nd half D-time

Details

UT 04:22 - Taking LBC Biases

UT 04:40 - Biases complete

UT 04:43 - Opening

UT 06:55 - Cirrus and clouds visible on all-sky image

UT 07:25 - Switching over to Science D-Time - first program is CV recovery of Saturnian Moon Greip

UT 07:46 - Running DOFPIA

UT 07:58 - Focus done -0.94" blue - 0.75" red

UT 08:00 - Running co-pointing - DIMM 0.85"

UT 08:03 - Copointing complete

UT 08:04 - Starting Greip05122.OB - Ran NSGUI to hijack preset

UT 08:30 - Completed Grep05122.OB script

UT 08:31 - Have noticed on chip 1 in LBC-R a "shadow" on the right edge of the chip. See IT#2659

UT 08:34 - Running up LUCI-1, need to reconnect TCS

UT 08:37 - Authorizing LUCI-1 to test NS Guiding with LUCI-1 on a comet

UT 08:43 - Taking 5 test darks

UT 08:58 - At low elevation - will make attempt

UT 09:14 - executing Comet_281P_2015.img script with NSIGUI hijack

UT 09:31 - Finished Comet test - moving to PDS 456

UT 09:33 - Authorizing LBC-R and switching to Binocular Mode

UT 09:50 - Collimated and sending preset from LUCI

UT 09:51 - Sending preset from LBC-R

UT 09:52 - Unwrapping Telescope

UT 09:55 - Resending LUCI-1 Presert, LBC-R being sent - Airmass 1.539

UT 10:04 - #10 acq image, #11 slit image

UT 10:09 - Starting spectroscopic script - airmass 1.577 (luci.20150522.0013-0017)

UT 10:12 - Started LBC-R OB

UT 10:48 - Had an LBC-R alarm on the OA computer, but nothing appeared on the GUI log or in lbccontrol's lbc.log

UT 11:01 - Alarm on Red again. From LBC.log:

2015/05/22 11:01:05.328000 N R CAMERA CCDCTRL integration completed
2015/05/22 11:01:05.843000 S R CAMERA CCDCTRL electronics temperature: 41.4 [Celsius]
2015/05/22 11:01:06.359000 S R CAMERA CCDCTRL original temperature reading value: 38096.708464
2015/05/22 11:01:06.359000 S R CAMERA CCDCTRL detectors temperature: 179.9 [Kelvin]

UT 11:10 - stopping LBC-R loop to focus

UT 11:31 - Stopped at 5th dither position for LUCI-1, authorizing LUCI-1 so can go to Telluriuc HIP86523

UT 11:39 - Telluric spectra started

UT 11:47 - Closing Enclosure

UT 11:59 - Taking HK200 zJ flat

UT 12:06 - Taking HK200 zJ Arc

-- %USERSIG{BarryRothberg - 2015-05-22}%

Comments

 
Topic revision: r2 - 22 May 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback